# Rice-like complexity lower bounds for Boolean and uniform automata networks

### v1

@@ 5 Arithmetics for uniform networks
In this section we study the intersection of pairs of number sequences, to the purpose of constructing a metareduction proving Theorem 1 for $q$-uniform automata networks (with a fixed alphabet of size $q$). We start by considering the restriction to Boolean alphabets, which is simpler. In Subsection 5.1, we extract regularities in the intersection of the two sets of integers $\{ak+b\mid k\in\mathbb{N}_{+}\}$ and $\{2^{n}\mid n\in\mathbb{N}\}$, respectively corresponding to the sizes of (counter-)model graphs produced by the pumping technique, and the sizes of admissible graphs as the dynamics of a Boolean AN (on configurations $X=\{0,1\}^{n}$). In Subsection 5.2, we immediately sketch how to derive Theorem 1 for $q=2$. In Subsection 5.3, we study the intersection of $\{ak+b\mid k\in\mathbb{N}_{+}\}$ and $\{q^{n}\mid n\in\mathbb{N}\}$, and again extract a geometric subsequence suitable for the reduction. Based on these arithmetical considerations, we will expose the proof of Theorem 1 for $q$-uniform networks (for all $q\geq 2$) in Section 6.
@@ 5.1 Geometric sequence for Boolean alphabets
  • the size of graphs obtained by pumping, i.e. from one base graph of size $b$ by gluing $k$ copies of graphs of size $a$ (taking into account the merged ports),
@@
If $b$ is odd and $a$ is even, then there is no solution, i.e. ${\mathcal{P}(a,b)}=\emptyset$.
@@
If $a\leq b$ we can write $b=am+b^{\prime}$ with $m\in\mathbb{N}_{+}$ and $0\leq b^{\prime}<a$ by Euclidean division. Observe that if $(K,N)$ is a solution for $(a,b)$, then $(K+m,N)$ is a solution for $(a,b^{\prime})$, because $2^{N}=aK+b=a(K+m)+b^{\prime}$. Conversely, if $(K,N)$ is a solution for $(a,b^{\prime})$, then $(K-m,N)$ is a solution for $(a,b)$, provided that $K-m>0$. The geometric sequences contained in ${\mathcal{P}(a,b)}$ and ${\mathcal{P}(a,b^{\prime})}$ are identical, up to a shift of the first term. Consequently Theorem 7 holds for $(a,b)$ if and only if it holds for $(a,b^{\prime})$, and we can restrict its study to the pairs $(a,b)$ with $a>b$.
@@ 5.2 Proof of Theorem 1 for $q=2$
For any $2$-non-trivial (resp. $2$-arborescent) MSO formula $\psi$, problem $\psi$-2-AN-dynamics (resp. $\psi$-2-NAN-dynamics) is ${\mathsf{NP}}$- or ${\mathsf{coNP}}$-hard.
@@
Consider a $2$-non-trivial (resp. $2$-arborescent) formula $\psi$. Without loss of generality, assume that the saturating graph $\Omega_{m}$ from Corollary 3 (resp. Proposition 2) is a counter-model of $\psi$, and let us build a reduction from UNSAT (otherwise consider $\neg\psi$ instead of $\psi$, and a reduction from SAT instead of UNSAT).
@@
From Proposition 5 (resp. Proposition 6) we obtain three $k$-graphs $\tilde{\Gamma}=\{\tilde{G}_{1},\tilde{G}_{2},\tilde{G}_{3}\}$, but their construction does not ensure that $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)$ is the dynamics of a $2$-uniform automata network for any $\ell\in\mathbb{N}$. The only obstacle for this is the size of $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)$, which is not guaranteed to be a power of $2$. From the construction, we only know that $\Delta^{\tilde{\Gamma}}(2\cdot 1\cdot 3)=\tilde{G}_{2}\oplus\tilde{G}_{1}\oplus\tilde{G}_{3}$ is a Boolean model of $\psi$ (from the $2$-non-triviality or $2$-arborescence). Hence, from $P_{1}(\tilde{G}_{1})\cap P_{2}(\tilde{G}_{1})\neq V(\tilde{G}_{1})$, we deduce the existence of $(a,b)\in\mathbb{N}_{+}\times\mathbb{N}$ with:
  • $b=|\tilde{G}_{2}\oplus\tilde{G}_{1}\oplus\tilde{G}_{3}|=|\tilde{G}_{2}|+|\tilde{G}_{1}|+|\tilde{G}_{3}|-2k$,
  • $a=|\tilde{G}_{1}|-k$,
@@
such that $|\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)|=a\ell+b$. When this size is a power of $2$, then $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)$ is the dynamics of a BAN suitable for the reduction. We have the initial solution $(1,\log_{2}(a+b))$ for $\ell=1$ (where $a+b$ is a power of two), therefore by Theorem 7, in $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)$ we have [^1] Euler’s totient function at $a\in\mathbb{N}_{+}$ is denoted $\varphi(a)$. graphs of size $(a+b)2^{\ell^{\prime}\varphi(a)}$ for all $\ell^{\prime}\in\mathbb{N}$, precisely where $\ell=\frac{(a+b)2^{\ell^{\prime}\varphi(a)}-b}{a}$ are integers (number of copies of $\tilde{G}_{1}$, which is $1$ for $\ell^{\prime}=0$).
@@
Before continuing, let us consider the fourth graph. Let $\tilde{G_{0}}=\Omega_{m}\sqcup\tilde{G_{1}}$. Let $\alpha$ be the least power of $2$ such that $G_{1}=\bigoplus_{i=1}^{\alpha}\tilde{G}_{1}$ verifies $|G_{1}|\geq|\tilde{G}_{0}|$, and let $G_{0}$ be the disjoint union of $\tilde{G}_{0}$ and $|G_{1}|-|\tilde{G}_{0}|$ fixed points, so that $|G_{0}|=|G_{1}|$. Finaly, let $G_{2}=\tilde{G}_{2}$ and $G_{3}=\tilde{G}_{3}$. The $k$-graphs $\Gamma=\{G_{0},G_{1},G_{2},G_{3}\}$ verify Conditions 1–3 of Proposition 4, as usual. Furthermore, $G_{1}$ is made of $\alpha$ copies of $\tilde{G}_{1}$.
@@
Given a propositional formula $S$ on $s$ variables, we need at least $2^{s}$ copies of $G_{0}$ or $G_{1}$ in order to implement the reduction from UNSAT, where for each valuation the circuit of the AN (resp. NAN) builds a copy of $G_{0}$ if $S$ is satisfied, and a copy of $G_{1}$ otherwise (see Section 3.2). Recall that $G_{1}$ is made of $\alpha$ copies of $\tilde{G}_{1}$ and $|G_{0}|=|G_{1}|$, that is, for $|w|=2^{s}$ we have $|\Delta^{\Gamma}(2\cdot w\cdot 3)|=a\alpha 2^{s}+b$. The goal is now to find a suitable padding by $L(s)$ copies of $\tilde{G}_{1}$, i.e. such that:
@@
is a power of $2$, with $|w|=2^{s}$, $G_{4}=\tilde{G}_{1}$ and $\Gamma^{\prime}=\{G_{0},G_{1},G_{2},G_{3},G_{4}\}$. Then the reduction is as usual: the graph $\Delta^{\Gamma^{\prime}}(2\cdot w\cdot 4^{L(s)}\cdot 3)$ is a model of $\psi$ if and only if $w$ does not contain letter 00 (in this case it has only copies of $\tilde{G}_{1}$). Remember that the only non-constant part is $S$ (and $s$).
@@
A solution with $L(s)\geq 0$ is $\ell^{\prime}=s+\log_{2}(\alpha)$, because the middle term is positive222$\ell^{\prime}=\left\lceil\frac{s+\log_{2}(\alpha)}{\varphi(a)}\right\rceil$ would be a smaller valid solution, but ceilings are not handy. For the same reason, $\alpha$ has been chosen to be a power of $2$.. Therefore, we can build a Boolean AN (resp. Boolean NAN) with
@@
$\alpha 2^{s}+L(s)=\frac{(a+b)\alpha^{\varphi(a)}2^{s\varphi(a)}-b}{a}$ copies of $\tilde{G}_{1}$
(some copies of $\tilde{G}_{1}$ form copies of $G_{1}$, or are replaced by a copy of $G_{0}$ when $S$ is satisfied). Recall that $a$, $b$, $\varphi(a)$ and $\alpha$ are constants depending only on $\psi$.
@@
  • the copy of $G_{2}$: from 00 to $|G_{2}|-1$,
  • one of the $2^{s}$ copies of $G_{0}$ or $G_{1}$ depending on $S$: from $|G_{2}|$ to $|G_{2}|+2^{s}(|G_{1}|-k)-1$, where $\beta=\lceil\log_{2}(\frac{x-|G_{2}|}{|G_{1}|-k})\rceil$ indicates the valuation from 00 to $12^{s}-$ 1, and $x-|G_{2}|-2^{\beta}(|G_{1}|-k)$ indicates the position within a copy of $G_{0}$ or $G_{1}$,
  • the copy of $G_{3}$ at the end of the gluing: from $|G_{2}|+2^{s}(|G_{1}|-k)$ to $|G_{2}|+2^{s}(|G_{1}|-k)+|G_{3}|-k-1$,
  • one of the $L(s)$ copies of $G_{4}=\tilde{G}_{1}$: from $|G_{2}|+2^{s}(|G_{1}|-k)+|G_{3}|-k$ to $12^{n}-$ 1.
@@ 5.3 Geometric sequence for $q$-uniform alphabets
Let us consider an arbitrary integer $q\geq 2$. We adapt the developments of Subsection 5.1, in order to produce $q$-uniform networks (of size $q^{n}$) via the pumping technique. We denote ++=$\mathbb{N}_{++}=\mathbb{N}\setminus\{0,1\}$, so that $q\in\mathbb{N}_{++}$. Unless specified explicitly differently, we consider $(a,b)\in\mathbb{N}_{+}\times\mathbb{N}$. Let
@@
If $b$ is not divisible by $q$ and $a$ is divisible by $q$, then either there is a unique solution $(K,N)\in\mathbb{N}^{2}$ with $N=0$, or there is no solution, i.e. ${\mathcal{P}_{q}(a,b)}=\emptyset$.
@@
If $a\leq b$ we can write $b=am+b^{\prime}$ with $m\in\mathbb{N}_{+}$ and $0\leq b^{\prime}<a$ by Euclidean division. Observe that if $(K,N)$ is a solution for $(a,b,q)$, then $(K+m,N)$ is a solution for $(a,b^{\prime},q)$, because $q^{N}=aK+b=a(K+m)+b^{\prime}$. Conversely, if $(K,N)$ is a solution for $(a,b^{\prime},q)$, then $(K-m,N)$ is a solution for $(a,b,q)$, provided that $K-m>0$. The geometric sequences contained in ${\mathcal{P}_{q}(a,b)}$ and ${\mathcal{P}_{q}(a,b^{\prime})}$ are identical, up to a shift of the first term. Consequently the first part of Theorem 10 holds for $(a,b,q)$ if and only if it holds for $(a,b^{\prime},q)$, and we can restrict its study to the triplets with $a>b$.
@@
By Lemma 13, it is suffisiant to prove the claim for $a>b$ and $a$ not divisible by $q$. Assume that ${\mathcal{P}_{q}(a,b)}\neq\emptyset$, i.e. there is a solution $(K,N)$ with $aK+b=q^{N}$. Assume furthermore that we have $(\mu,\kappa)\in\mathbb{N}_{+}\times\mathbb{N}_{+}$ such that $a\kappa+b=bq^{\mu}$.
@@
By Lemmas 11 and 15 (recall that $\kappa=0$ implies $b=0$). ∎
@@
For any $(a,q)\in\mathbb{N}_{+}\times\mathbb{N}_{++}$, define the coprime power of $q$ for $a$ as the smallest integer $\eta$ such that when writing $a=a^{\prime}\gcd(a,q^{\eta})$, $a^{\prime}$ is coprime with $q$.
@@
Finaly, we prove that $\eta$ is the smallest integer with such a property. For the sake of a contradiction, assume that there exists $\mu<\eta$ with the same property. Consequently, there exists at least one prime number $p\in\mathcal{P}$ such that $v_{p}(q)\neq 0$ and $\min(v_{p}(a),\mu v_{p}(q))<\min(v_{p}(a),\eta v_{p}(q))=v_{p}(a)$. Hence, $a=a^{\prime\prime}p\gcd(a,q^{\mu})$ for some $a^{\prime\prime}\in\mathbb{N}_{+}$ and $a^{\prime\prime}p$ is not coprime with $q$, which is a contradiction. ∎
@@
If $q$ is prime, then we can assume that $a$ is coprime with $q$ (otherwise we employ Lemma 13, because $a$ is divisible by $q$). As a consequence, $a^{\prime}=a$ and $\eta=0$ in Lemma  17, so the sufficient condition is verified for any $b\in\mathbb{N}_{+}$, and by Theorem 10 there is a geometric sequence of integers in ${\mathcal{P}_{q}(a,b)}$.
@@
Combining Lemmas 17 and 18, we obtain an equivalent condition.

### v2

@@ 5 Arithmetics for uniform networks
In this section we study the intersection of pairs of number sequences, to the purpose of constructing a metareduction proving Theorem 1 for $q$-uniform automata networks (with a fixed alphabet of size $q$). We start by considering the restriction to Boolean alphabets, which is simpler. In Subsection 5.1, we extract regularities in the intersection of the two sets of integers $\{ak+b\mid k\in\mathbb{N}_{+}\}$ and $\{2^{n}\mid n\in\mathbb{N}\}$, respectively corresponding to the sizes of (counter-)model graphs produced by the pumping technique, and the sizes of admissible graphs as the dynamics of a Boolean automata network (on configurations $X=\{0,1\}^{n}$). In Subsection 5.2, we immediately sketch how to derive Theorem 1 for $q=2$. In Subsection 5.3, we study the intersection of $\{ak+b\mid k\in\mathbb{N}_{+}\}$ and $\{q^{n}\mid n\in\mathbb{N}\}$, and again extract a geometric subsequence suitable for the reduction. Based on these arithmetical considerations, we will expose the proof of Theorem 1 for $q$-uniform networks (for all $q\geq 2$) in Section 6.
@@ 5.1 Geometric sequence for Boolean alphabets
  • the size of graphs obtained by pumping, i.e. from one base graph of size $b$ by gluing $k$ copies of graphs of size $a$,
@@
If $b$ is odd and $a$ is even, then there is no solution, i.e. $\mathcal{P}(a,b)$$\emptyset$.
@@
If $a\leq b$ we can write $b=am+b^{\prime}$ with $m\in\mathbb{N}_{+}$ and $0\leq b^{\prime}<a$ by Euclidean division. Observe that if $(K,N)$ is a solution for $(a,b)$, then $(K+m,N)$ is a solution for $(a,b^{\prime})$, because $2^{N}=aK+b=a(K+m)+b^{\prime}$. Conversely, if $(K,N)$ is a solution for $(a,b^{\prime})$, then (Km,N)(K-m,N) is a solution for $(a,b)$, provided that Km>0K-m>0. The geometric sequences contained in $\mathcal{P}(a,b)$} and $\mathcal{P}(a,b^{\prime})$} are identical, up to a shift of the first term. Consequently Theorem 10 holds for $(a,b)$ if and only if it holds for $(a,b^{\prime})$, and we can restrict its study to the pairs $(a,b)$ with $a>b$.
@@ 5.2 Proof of Theorem 1 for $q=2$
For any $2$-non-trivial (resp. $2$-cw-non-trivial) MSO formula $\psi$, problem $\psi$-2-AN-dynamics (resp. $\psi$-2-NAN-dynamics) is $\mathsf{NP}$}- or $\mathsf{coNP}$}-hard.
The deterministic case is handled as in Section 4, to ensure that all the graphs have out-degree $1$. No other detail is needed, the main consideration here is on graph sizes.
@@
Consider a $2$-non-trivial (resp. $2$-cw-non-trivial) formula $\psi$. Without loss of generality, assume that the saturating graph $\Omega_{m}$ from Corollary 4 (resp. Proposition 3) is a counter-model of $\psi$, and let us build a reduction from UNSAT (otherwise consider $\neg\psi$ instead of $\psi$, and a reduction from SAT instead of UNSAT).
@@
From Proposition 5 (resp. Proposition 9) we obtain three clique-decompositions $\tilde{\Gamma}=\{\tilde{\mathcal{C}}_{1},\tilde{\mathcal{C}}_{2},\tilde{\mathcal{C}}_{3}\}$ of cliquewidth at most $k$, but their construction does not ensure that the graph generated by $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)$ is the dynamics of a $2$-uniform automata network for any $\ell\in\mathbb{N}$. The only obstacle for this is the size of $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)$, which is not guaranteed to be a power of $2$. From the construction, we only know that $\Delta^{\tilde{\Gamma}}(2\cdot 1\cdot 3)=\tilde{\mathcal{C}}_{2}\oplus\tilde{\mathcal{C}}_{1}\oplus\tilde{\mathcal{C}}_{3}$ is a Boolean model of $\psi$ (from the $2$-non-triviality or $2$-cw-non-triviality). Hence, we deduce the existence of $(a,b)\in\mathbb{N}_{+}\times\mathbb{N}$ with:
  • $b=|\tilde{\mathcal{C}}_{2}\oplus\tilde{\mathcal{C}}_{1}\oplus\tilde{\mathcal{C}}_{3}|=|\tilde{\mathcal{C}}_{2}|+|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{3}|$,
  • $a=|\tilde{\mathcal{C}}_{1}|$,
@@
such that $|\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)|=a\ell+b$. When this size is a power of $2$, then $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)$ is the dynamics of a BAN suitable for the reduction. We have the initial solution $(1,\log_{2}(a+b))$ for $\ell=1$ (where $a+b$ is a power of two), therefore by Theorem 10, in $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)$ we have graphs of size $(a+b)2^{\ell^{\prime}\varphi(a)}$ for all $\ell^{\prime}\in\mathbb{N}$, precisely where $\ell=\frac{(a+b)2^{\ell^{\prime}\varphi(a)}-b}{a}$ are integers (number of copies of $\tilde{\mathcal{C}}_{1}$, which is $1$ for $\ell^{\prime}=0$).
@@
Before continuing, let us consider the fourth clique-decomposition. Let $\tilde{\mathcal{C}_{0}}=\mathcal{C}_{m}\sqcup\tilde{\mathcal{C}_{1}}$, where $\mathcal{C}_{m}$ is a clique-decomposition of $\Omega_{m}$. Let $\alpha$ be the least power of $2$ such that $\mathcal{C}_{1}=\bigoplus_{i=1}^{\alpha}\tilde{\mathcal{C}}_{1}$ verifies $|\mathcal{C}_{1}|\geq|\tilde{\mathcal{C}}_{0}|$, and let $\mathcal{C}_{0}$ be the disjoint union of $\tilde{\mathcal{C}}_{0}$ and $|\mathcal{C}_{1}|-|\tilde{\mathcal{C}}_{0}|$ constant nodes, so that $|\mathcal{C}_{0}|=|\mathcal{C}_{1}|$. Finally, let $\mathcal{C}_{2}=\tilde{\mathcal{C}}_{2}$ and $\mathcal{C}_{3}=\tilde{\mathcal{C}}_{3}$. The clique-decompositions $\Gamma=\{\mathcal{C}_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}$ of cliquewidth at most $k$ verify Conditions (i)–(iii) of Proposition 7. Furthermore, $\mathcal{C}_{1}$ is made of $\alpha$ copies of $\tilde{\mathcal{C}}_{1}$.
@@
Given a propositional formula $S$ on $s$ variables, we need at least $2^{s}$ copies of $\mathcal{C}_{0}$ or $\mathcal{C}_{1}$ in order to implement the reduction from UNSAT, where for each valuation the circuit of the AN (resp. NAN) produces a copy of $\mathcal{C}_{0}$ if $S$ is satisfied, and a copy of $\mathcal{C}_{1}$ otherwise (see Subsection 3.3). Recall that $\mathcal{C}_{1}$ is made of $\alpha$ copies of $\tilde{\mathcal{C}}_{1}$ and $|\mathcal{C}_{0}|=|\mathcal{C}_{1}|$, that is, for $|w|=2^{s}$ we have $|\Delta^{\Gamma}(2\cdot w\cdot 3)|=a\alpha 2^{s}+b$. The goal is now to find a suitable padding by $L(s)$ copies of $\tilde{\mathcal{C}}_{1}$, i.e. such that:
@@
is a power of $2$, with $|w|=2^{s}$, $\mathcal{C}_{4}=\tilde{\mathcal{C}}_{1}$ and $\Gamma^{\prime}=\{\mathcal{C}_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3},\mathcal{C}_{4}\}$. Then the reduction is as usual: the graph $\Delta^{\Gamma^{\prime}}(2\cdot w\cdot 4^{L(s)}\cdot 3)$ is a model of $\psi$ if and only if $w$ does not contain letter 0 (in this case it has only copies of $\tilde{\mathcal{C}}_{1}$). Remember that the only non-constant part is $S$ (and $s$).
@@
[…]
@@
$\alpha 2^{s}+L(s)=\frac{(a+b)\alpha^{\varphi(a)}2^{s\varphi(a)}-b}{a}$ copies of $\tilde{\mathcal{C}}_{1}$
(some copies of $\tilde{\mathcal{C}}_{1}$ form copies of $\mathcal{C}_{1}$, or are replaced by a copy of $\mathcal{C}_{0}$ when $S$ is satisfied). Recall that $a$, $b$, $\varphi(a)$ and $\alpha$ are constants depending only on $\psi$.
@@
  • the copy of $\mathcal{C}_{2}$: from 0 to $|\mathcal{C}_{2}|-1$,
  • one of the $2^{s}$ copies of $\mathcal{C}_{0}$ or $\mathcal{C}_{1}$ depending on $S$: from $|\mathcal{C}_{2}|$ to $|\mathcal{C}_{2}|+2^{s}|\mathcal{C}_{1}|-1$, where $\beta=\lfloor\frac{x-|\mathcal{C}_{2}|}{|\mathcal{C}_{1}|}\rfloor$ indicates the valuation from 0 to $2^{s}-1$, and $x-|\mathcal{C}_{2}|-\beta|\mathcal{C}_{1}|$ indicates the position within a copy of $\mathcal{C}_{0}$ or $\mathcal{C}_{1}$,
  • the copy of $\mathcal{C}_{3}$ at the end of the gluing: from $|\mathcal{C}_{2}|+2^{s}|\mathcal{C}_{1}|$ to $|\mathcal{C}_{2}|+2^{s}|\mathcal{C}_{1}|+|\mathcal{C}_{3}|-1$,
  • one of the $L(s)$ copies of $\mathcal{C}_{4}=\tilde{\mathcal{C}}_{1}$: from $|\mathcal{C}_{2}|+2^{s}|\mathcal{C}_{1}|+|\mathcal{C}_{3}|$ to $2^{n}-1$.
@@ 5.3 Geometric sequence for $q$-uniform alphabets
Let us consider an arbitrary integer $q\geq 2$. We adapt the developments of Subsection 5.1, in order to produce $q$-uniform networks (of size $q^{n}$) via the pumping technique. We denote $\mathbb{N}_{++}=\mathbb{N}\setminus\{0,1\}$, so that $q\in\mathbb{N}_{++}$. Unless specified explicitly differently, we consider $(a,b)\in\mathbb{N}_{+}\times\mathbb{N}$. Let
@@
If $b$ is not divisible by $q$ and $a$ is divisible by $q$, then either there is a unique solution $(K,N)\in\mathbb{N}^{2}$ with $N=0$, or there is no solution, i.e. $\mathcal{P}_{q}(a,b)$$\emptyset$.
@@
If $a\leq b$ we can write $b=am+b^{\prime}$ with $m\in\mathbb{N}_{+}$ and $0\leq b^{\prime}<a$ by Euclidean division. Observe that if $(K,N)$ is a solution for $(a,b,q)$, then $(K+m,N)$ is a solution for $(a,b^{\prime},q)$, because $q^{N}=aK+b=a(K+m)+b^{\prime}$. Conversely, if $(K,N)$ is a solution for $(a,b^{\prime},q)$, then (Km,N)(K-m,N) is a solution for $(a,b,q)$, provided that Km>0K-m>0. The geometric sequences contained in $\mathcal{P}_{q}(a,b)$} and $\mathcal{P}_{q}(a,b^{\prime})$} are identical, up to a shift of the first term. Consequently the first part of Theorem 13 holds for $(a,b,q)$ if and only if it holds for $(a,b^{\prime},q)$, and we can restrict its study to the triplets with $a>b$.
@@
By Lemma 16, it is sufficient to prove the claim for $a>b$ and $a$ not divisible by $q$. Assume that $\mathcal{P}_{q}(a,b)$$\neq\emptyset$, i.e. there is a solution $(K,N)$ with $aK+b=q^{N}$. Assume furthermore that we have $(\mu,\kappa)\in\mathbb{N}_{+}\times\mathbb{N}_{+}$ such that $a\kappa+b=bq^{\mu}$.
@@
By Lemmas 14 and 18 (recall that $\kappa=0$ implies $b=0$). ∎
@@
For any $(a,q)\in\mathbb{N}_{+}\times\mathbb{N}_{++}$, define the coprime power of $q$ for $a$ as the smallest integer $\eta$ such that when writing $a^{\prime}\gcd(a,q^{\eta}), aa^{\prime}$ is coprime with $q$.
@@
Finally, we prove that $\eta$ is the smallest integer with such a property. For the sake of a contradiction, assume that there exists $\mu<\eta$ with the same property. Consequently, there exists at least one prime number $p\in\mathcal{P}$ such that $v_{p}(q)\neq 0$ and $\min(v_{p}(a),\mu v_{p}(q))<\min(v_{p}(a),\eta v_{p}(q))=v_{p}(a)$. Hence, $a=a^{\prime\prime}p\gcd(a,q^{\mu})$ for some $a^{\prime\prime}\in\mathbb{N}_{+}$ and $a^{\prime\prime}p$ is not coprime with $q$, which is a contradiction. ∎
@@
If $q$ is prime, then we can assume that $a$ is coprime with $q$ (otherwise we employ Lemma 16, because $a$ is divisible by $q$). As a consequence, $a^{\prime}=a$ and $\eta=0$ in Lemma 20, so the sufficient condition is verified for any $b\in\mathbb{N}_{+}$, and by Theorem 13 there is a geometric sequence of integers in $\mathcal{P}_{q}(a,b)$}.
@@
Combining Lemmas 20 and 21, we obtain an equivalent condition.